OpenStudy (anonymous):

solve the system using elimination 6q - r + 2s = 8 2q + 3r - 5 = -9 4q + 2r + 5s = 1 Steps would be appreciated u w u

OpenStudy (tkhunny):

1) Fix the second equation. 2) Pick a variable and solve for it in any one equation. 3) Substitute for the variable the solved-for expression in the other equations. 4) Repeat 2) and 3) as necessary. That's about it.

OpenStudy (anonymous):

Ok, I'm really, really lost... How would I complete those steps? Sorry :/

OpenStudy (tkhunny):

You must do #1. No one can help you with that.

OpenStudy (anonymous):

Real helpful, dude. My problem is that I've no idea how to do number one.

OpenStudy (texaschic101):

he has to do it by elimination, not substitution tkhunny

OpenStudy (anonymous):

Or any of the problems for that matter :/

OpenStudy (tkhunny):

Seriously? You can't fix the equation that is written incorrectly?

OpenStudy (anonymous):

I don't know what's wrong with it!!!!

OpenStudy (tkhunny):

Oh, right. Misread that. Anyway, you still have to do #1.

OpenStudy (texaschic101):

you forgot the s

OpenStudy (tkhunny):

The "s" might be missing.

OpenStudy (texaschic101):

is equation 2 supposed to be : 2q - 3r - 5s = -9 ??

OpenStudy (anonymous):

Yeah. Sorry, I copied it incorrectly.

OpenStudy (tkhunny):

There you go. So we see that you can do step #1. Now, listen to Texas.

OpenStudy (tkhunny):

1) 6q - r + 2s = 8 2) 2q + 3r - 5s = -9 3) 4q + 2r + 5s = 1 The idea is to ELIMINATE things. It's kind of magic. I numbered the equations for simplicity of communication. Calculate 1) - 3*2) Calculate 3) - 2*2) It will be fun!